40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Jul 16 16:00:01 2019, gautam, Update, CDS, c1iscaux Supermicro setup IMG_7769.JPG
    Reply  Wed Jul 17 21:22:41 2019, gautam, Update, CDS, CM board Latch Enable subtlety 
       Reply  Sun Jul 21 12:55:38 2019, gautam, Update, CDS, CM board Latch Enable test script LatchLogic.pdfLatchLogicTest.png
          Reply  Fri Aug 9 08:59:04 2019, gautam, Update, CDS, Prep for install of c1iscaux 
             Reply  Sun Aug 11 11:47:42 2019, gautam, Update, CDS, Bench test of c1iscaux 
                Reply  Mon Aug 12 17:36:04 2019, gautam, Update, CDS, More bench test of c1iscaux iscauxCheclist.pdf
                   Reply  Mon Aug 12 21:25:19 2019, Koji, Update, CDS, More bench test of c1iscaux 
                   Reply  Tue Aug 13 08:07:09 2019, gautam, Update, CDS, P1--->P2 
                      Reply  Tue Aug 13 14:36:17 2019, gautam, Update, CDS, P1--->P2 
                   Reply  Fri Aug 16 16:40:04 2019, gautam, Update, CDS, 1Y3 work newLook1Y3.JPGIMG_7803.JPGc1lsc_crashed.png
                      Reply  Sat Aug 17 16:49:23 2019, gautam, Update, CDS, More 1Y3 work Screen_Shot_2019-08-17_at_3.00.57_PM.pngScreen_Shot_2019-08-17_at_3.12.23_PM.pngIMG_7804.JPGScreenshot_from_2019-08-17_17-04-47.png
                         Reply  Mon Aug 19 14:36:21 2019, gautam, Update, CDS, c1iscaux remaining work caseForSmallerFootprint.pdf
                            Reply  Tue Aug 20 19:05:24 2019, Koji, Update, CDS, MC1 (and MC3) troubleshoot Screenshot_from_2019-08-20_17-26-01.pngScreenshot_from_2019-08-20_17-43-03.png
                               Reply  Thu Aug 22 12:54:06 2019, Koji, Update, CDS, MC1 glitch removed (for now) and IMC locking recovered 7x
                                  Reply  Thu Aug 22 20:56:51 2019, Koji, Update, CDS, MC1 glitch removed (for now) and IMC locking recovered 
                            Reply  Fri Aug 23 18:46:17 2019, Jon, Update, CDS, c1iscaux remaining work 
                               Reply  Sun Aug 25 14:18:08 2019, gautam, Update, CDS, c1iscaux remaining work Screen_Shot_2019-08-25_at_10.38.37_PM.png
    Reply  Thu Jul 18 10:46:04 2019, gautam, Update, CDS, Database files made 
       Reply  Thu Jul 18 19:58:56 2019, gautam, Update, CDS, Work on Acromag chassis IMG_7771.JPGIMG_7770.JPGIMG_7772.JPG
          Reply  Fri Jul 19 19:44:03 2019, gautam, Update, CDS, Database file test Whitening.png
Message ID: 14845     Entry time: Tue Aug 13 14:36:17 2019     In reply to: 14844
Author: gautam 
Type: Update 
Category: CDS 
Subject: P1--->P2 

As it turns out, only one extra shroud needed to be installed - I did this and migrated the cables for the 4 whitening boards from the P1 to P2 connectors. So until the new Acromag box is installed, we have no control over the whitening gains (slow channels), but do still have control over the whitening filter enable/disable (controlled by fast BIO). I am thinking about the easiest way to test the latter - I think the ambient PD dark noise level is too low to be seen above ADC noise even with the whitening enabled, and setting up drive signals to individual channels is too painful - maybe with +45dB of whitening gain, the (z,p) whitening filter shape can be seen with just PD/demod chain electroncis noise.

Quote:

This morning, I wanted to move the existing cables going to the P1 connectors of the iLIGO whitening boards to the P2 connector, to test the modifications made to allow whitening stage switching. Unfortunately, I found that the shrouds werent installed. Where can I find these?

ELOG V3.1.3-